Given that the quantity of ozone in the stratosphere is equivalent to a 3.0-mm-thick layer of ozone on Earth at STP, calculate the number of ozone molecules in the stratosphere and their mass in kilograms. (Hint: The radius of Earth is 6371 km and the surface area of a sphere is 4 pi r^2, where r is the radius).


1
Expert's answer
2021-01-13T13:54:07-0500

The radius of earth (r) = 6371 km

Surface area of a sphere (S.A) = 4πr² = "4× \dfrac{22}7 × (6371 ×10^3)^2" = 5.1 × 10¹⁴ m²


Volume of the atmosphere volume of atmosphere = S.A × thickness = 5.1 × 10¹⁴ × 3 × 10-³ = 15.3 × 10¹¹ m³


Since 1m³ = 1000L

15.3 × 10¹¹ m³= 15.3 × 10¹⁴L



Since 22.4 L contains 6.023 × 10²³ Ozone molecules,

Thus, 15.3 × 10¹⁴ L = "\dfrac{6.023×10²³ × 15.3 × 10¹⁴L}{22.4}" = 4.11 × 10³⁷ molecules of Ozone



1 mole of Ozone (O3) = 48g (16×3g)

6.023 × 10²³ molecules of Ozone = 48g

4.11 × 10³⁷ molecules of Ozone = "\dfrac{4.11 × 10³⁷ × 48}{6.023× 10²³}" = 3.275 × 10¹⁵g = 3.275 × 10¹² kg



Therefore, the number of ozone molecules in their stratosphere and their mass in kilograms is 4.11 × 10³⁷ molecules and 3.275 × 10¹² kg respectively.